OpenStudy (anonymous):

Susan begins counting backward from 1298 by 4's, saying one number every 5 seconds. At the same time, Jim begins counting forward from 171 by 3's, saying one number every 5 seconds. What number will they both say at the same time? (A) 640 (B) 644 (C) 648 (D) 650 (E) 654

OpenStudy (anonymous):

@kelliegirl33 @kayla_b_stephens @nincompoop

OpenStudy (anonymous):

the questions and answer are on this exam sampler

OpenStudy (anonymous):

I go to Hunter College High School, but I wanted to practice math.

OpenStudy (anonymous):

oh okay well you can use it to check your answer then

OpenStudy (anonymous):

Well I wanted to know the explanation to approach each answer.

OpenStudy (anonymous):

i was trying to find one

OpenStudy (anonymous):

well...the answer is 654, but I am not sure how to do it. Kayla gave the answer sheet with that question on it...sorry I can't help

OpenStudy (anonymous):

i just did something that can give the answer

OpenStudy (anonymous):

what did you do ?

OpenStudy (anonymous):

1298/4 =324.5*2 +5= 654

OpenStudy (anonymous):

this is the only way i could get 654

OpenStudy (anonymous):

@myininaya ...can you explain how to do this please

OpenStudy (anonymous):

@e.mccormick ...can you help

OpenStudy (anonymous):

somebody will be able to explain this